### v1

# Deep Learning-Assisted Weak Beam Identification in Dark-Field X-ray Microscopy

Abstract
Dislocations control the mechanical behavior of crystalline materials, yet their quantitative characterization in bulk has remained elusive. Transmission Electron Microscopy provides atomic-scale resolution but is restricted to thin foils, limiting relevance to structural performance. Dark-field X-ray microscopy (DFXM) has recently opened access to three-dimensional, non-destructive imaging of dislocations in macroscopic crystals. A critical bottleneck, however, is the reliable identification of weak- versus strong-beam conditions. Weak-beam imaging enhances dislocation contrast, while strong-beam conditions are dominated by multiple scattering and obscure interpretation. Current practice depends on manual classification by specialists, which is subjective, slow, and incompatible with the scale of modern experiments. Here, we introduce a deep learning framework that automates this task using a lightweight convolutional neural network trained on small, hand-labeled datasets. By enabling robust, rapid, and scalable identification of imaging conditions, this approach transforms DFXM into a high-throughput tool, unlocking statistically significant studies of dislocation dynamics in bulk materials.
keywords:
Machine Learning, Diffraction Imaging, Dislocations, Dark Field MicroscopyIntroduction
Dislocations are fundamental to understanding both the mechanical and functional behavior of crystalline materials. As the primary carriers of plastic deformation, they mediate irreversible atomic motion within the lattice, dictating how materials yield, harden, and ultimately fail under applied stress [1, 2, 3]. Their role extends well beyond mechanical response: dislocations also shape key functional properties, including electrical conductivity, magnetic behavior, and atomic diffusion [4, 5, 6, 7]. Central to most experimental methods for probing dislocations is their angular sensitivity. In an ideal crystal, lattice planes are perfectly periodic, giving rise to sharp diffraction of incident X-rays or electrons at well-defined angles. Dislocations perturb this periodicity through their long-range strain fields, locally tilting lattice planes by fractions of a degree and shifting the diffraction condition by only milliradians. Owing to the high selectivity of diffraction, such subtle distortions generate measurable contrast, enabling direct imaging of dislocations.
Visualizing and characterizing dislocations across relevant length scales is essential for uncovering the mechanisms that govern deformation. Transmission electron microscopy (TEM) has played a foundational role in this effort, providing near-atomic resolution images that reveal dislocation core structures and interactions [8, 9]. However, the requirement for electron-transparent samples, typically limited to thicknesses below a few hundred nanometers, restricts TEM’s ability to probe three-dimensional dislocation networks within bulk crystals, especially those with low dislocation densities such as annealed single crystals. Additionally, for systems like semiconductors and ceramics, where dislocation densities are inherently low, the limited field of view in TEM poses a challenge for capturing statistically significant defect populations.
To overcome these limitations, dark-field X-ray microscopy (DFXM) has emerged as a powerful, non-destructive method for imaging deeply embedded dislocations under realistic conditions [10, 11, 12, 13]. Using high-energy synchrotron X-rays and objective-based imaging in Bragg diffraction geometry, DFXM enables real-space visualization of lattice distortions with sub-micrometer spatial resolution and angular sensitivity on the order of radians. Conceptually similar to dark-field TEM but with greater penetration depth and tunable magnification, DFXM achieves resolution limits around 50 nm. Notably, Jakobsen et al. [12] demonstrated the use of weak-beam (WB) contrast in DFXM, adapting a concept long used in TEM [14], to selectively image regions with high lattice distortion by recording at the tails of the rocking curve. This approach highlights dislocation lines with high specificity and clarity, providing detailed 2D projections of dislocation structures.
By mapping the local lattice orientation variations and elastic strain around a given diffraction vector, DFXM acquires spatially-resolved high-dimensional (up to 6D in static measurements) datasets that capture long-range strain fields and the spatial arrangement of dislocations in three dimensions. Datasets from subsequent virtual sections can be stacked to reconstruct volumetric images of defects, enabling in situ studies of dislocation network evolution under mechanical loading [16, 17, 18]. However, DFXM data analysis remains data intensive and requires significant supervision. In particular, this has become more prominent with the advent of extremely brilliant synchrotron sources (EBS) [19, 20] which allow mapping an entire grain within minutes [21].
Here we address a specific problem relevant for samples of high crystal quality, e.g. samples where individual dislocations within dislocation networks such as coherent small angle grain boundaries can be resolved. In such samples, strong multiple scattering occurs, and the dynamical theory of x-ray diffraction has to be applied in order to simulate the topographs observed near the maximum of the rocking curve [22, 23]. In the tails of the rocking curve, however, multiple scattering is negligible and the simpler kinematic theory of x-ray diffraction can be applied. This is called the WB condition. Therefore, the typical approach to analyzing RCI or DFXM data is to first segment the scanning data into weak- and strong-beam (SB) images, and them analyze and interpret the weak-beam images. The aim of this work is to provide a Machine Learning (ML) workflow for the automatic identification of weak-beam images in DFXM rocking curves. In subsequent data analysis steps, these weak beam images can the be used for the automatic identification of dislocations and other localized crystal defects.
Conventional rocking curves analysis methods rely on manual interpretation or unsupervised approaches such as principal component analysis [24] or Gram-Schmidt orthogonalization [25] to identify the weak-beam condition and visualize strain fields near defect cores. However, these methods are frequently time-consuming and lack generalizability across different datasets and material systems. As a result, there is a pressing need for more automated, robust, and flexible analysis pipelines.
Machine Learning (ML) algorithms can offer a promising solution for the above-mentioned challenges. In recent years, ML methods have been increasingly applied to diverse materials science problems [26, 27, 28, 29, 30]. The success of deep learning in computer vision has inspired materials scientists to adopt these techniques for tasks such as automatic classification and segmentation of microstructural features. ML approaches have shown promise in uncovering relationships between microscopic structures and material properties, supporting both theoretical studies and simulations. For example, Zhang et al. [31] used convolutional neural networks (CNNs) to extract dislocation microstructures from electron backscatter diffraction (EBSD) data, while Yassar et al. [32] employed artificial neural networks (ANNs) to predict the flow stress of micropillars using dislocation density and model size as input features. Similarly, Tao et al. [33] applied various ML models to predict micropillar compression behavior based on discrete dislocation dynamics (DDD) simulations, and Hiemer et al. [34] used Kernel Ridge Regression (KRR) to link plasticity with dislocation characteristics, analyzing the effects of strain rate, dislocation density, and strengthening mechanisms in FCC metals.
Despite their potential, deep neural networks often require large numbers of parameters and significant computational resources. To address this, lightweight network architectures [35, 36] have been developed to maintain accuracy while reducing model complexity and computational cost. Such models are well-suited for applications with limited computational resources. In the context of DFXM, the WB condition is not typically normally distributed and can appear irregularly within individual rocking curve frames. Moreover, not all data exhibit clearly separable weak- and strong-beam features, and user interpretation often plays a role in manual analysis. To address these challenges, we developed a patch-based ML workflow that reduces DFXM dataset size while automating the classification of diffraction contrast. This approach enables consistent identification of WB and SB conditions, streamlining defect analysis and reducing subjective bias. Our method offers a faster and more reliable route for analyzing DFXM dislocation data, and can be extended to related techniques that rely on similar contrast mechanisms, such as X-ray topography [37] and transmission electron microscopy [38].
Results
Deep learning model
Herein, we aim to develop automated methods for reducing and analyzing data from three-dimensional dislocation imaging using DFXM. A typical setup is illustrated in Figure 1A, where a single crystal or grain of interest is aligned through the objective with a given reflection. An incident X-ray beam illuminates the sample, and the diffracted beam is selected at the Bragg angle using a compound refractive lens (CRL) system and directed to a 2D detector. For typical 3D scans, a line-focused incident beam is used, and the sample is translated along the laboratory -axis to sequentially probe 2D virtual slices through the volume. At each -layer, the sample is rotated around an axis perpendicular to the incident beam, typically the -axis [39], to collect a rocking curve, with DFXM images acquired at each angular step, as shown in Figure 1B. Traditionally, WB images, those located at the tails of the rocking curve, are manually identified to enhance dislocation contrast. Accurate and reproducible reconstruction of the three-dimensional dislocation structure requires robust identification of appropriate weak-beam conditions, motivating the need for automated analysis tools.
To preserve all relevant information and enable reliable 3D reconstruction, we propose a patch-based approach for the localized identification of weak-beam conditions in each rocking layer. This method allows for small-patch-wise classification of beam conditions, a critical step for stacking and aligning the rocking layers. Ground truth labels for training were manually generated by sampling representative frames from the dataset and annotating them according to their corresponding beam condition.
For our the classification task of interest, we utilize a lightweight convolutional neural network (LCNN) based on a depthwise separable architecture [40, 41]. As depicted in Figure 2, the model operates on relatively small image patches pixels extracted from raw DFXM data.
The subdivision of full-sized raw images into smaller patches offers several advantages that enhance both the training efficiency and robustness of the model. By operating on smaller datasets, the LCNN reduces the number of variables, thereby improving computational efficiency. Additionally, in samples with structural complexity, such as strong curvature or subgrain boundaries, different regions of the image may exhibit the weak-beam condition at varying orientations. Using patches that are comparable in size to the characteristic features helps to avoid inconsistencies across the image. This localized approach also enables more effective use of labeled data; a single hand-labeled full-sized image can yield many smaller training patches (up to 1024 patches), significantly increasing the volume of training data without additional labeling effort. Once the model processes each patch, the segmented outputs can be reassembled into full-sized images for interpretation.
Each patch then undergoes a sequence of depthwise separable convolutional layers, beginning with 5×5 depthwise convolutions that extract spatial features within individual channels, followed by 1×1 pointwise convolutions that integrate information across channels. These features are projected into a 32-dimensional latent space, with non-linearities introduced via rectified linear unit (ReLU) activations to enhance representational capacity. The sequence is repeated multiple times to capture complex image characteristics, and the final representation is passed to a fully connected layer that classifies each patch as exhibiting either strong-beam or weak-beam conditions. To improve generalization and prevent overfitting, dropout regularization [42] with a probability of 0.15 is applied during training.
Model evaluation
The data used in this study was derived from a previously published and publicly available DFXM dataset [13]. Each rocking step frame had a resolution of 2160×2560 pixels, which was divided into non-overlapping 64×64 patches. Labeling was performed manually at the image level, meaning that entire images were classified as either SB or WB conditions. Based on this, 6 WB images and 6 SB images were randomly selected, resulting in 6144 patches for each beam condition. The ground truth dataset was split into 80% for training and 20% for testing. All patches were log-scaled to enhance feature contrast and were passed through the neural network in mini-batches during training to ensure efficient learning and generalization.
The model is trained for 50 epochs, converging around epoch 30, as shown in Figure 3. Both training and testing losses decrease steadily, with no significant divergence, indicating effective generalization. The use of dropout further enhances model robustness by deactivating a subset of neurons during training epochs. Classification performance is evaluated using accuracy, defined as the proportion of correctly classified samples relative to the total dataset size:
where TP and TN denote true positives and true negatives, respectively, and FP and FN represent false positives and false negatives. Additionally, the final epoch confusion matrix provides additional insight, showing 272 correctly classified SB patches and 67 correctly classified WB patches, with only 28 SB and 1 WB instances misclassified.
To benchmark our LCNN and to see how much we gained in terms of accuracy versus the computation cost, we compare with ResNet18 [43] and VGG16 [44] the other lightweight variants of the Resnet and VGG families. Table 1 summarizes their respective performance in terms of parameter count, training time, and accuracy.
| Model | Trainable Parameters | Training Time (S) | Accuracy (%) | |||
| LCNN | 879 | K | 23 | .87 | 92 | .70 |
| ResNet18 | 11 | M | 75 | .07 | 97 | .13 |
| VGG16 | 134 | M | 208 | .09 | 89 | .10 |
While ResNet18 achieves the highest classification accuracy, our LCNN provides a solid balance between performance and efficiency. Due to its lightweight architecture, it has significantly fewer parameters and trains faster, making it well-suited for real-time applications and environments with limited computational resources. Additionally, the patch size impacts both training/inference time and accuracy, therefore it is important for users to choose a patch size appropriate to their dataset. Based on our experiments, a patch size of pixels offers the best trade-off between accuracy and computational cost.
Discussion
After training, we proceed by assigning our LCNN to label the entire rocking curves by identifying WB regions, which are particularly of interest for dislocation mapping. To enhance defect visibility, frames corresponding to WB-labeled patches are extracted and combined into an integrated image. This approach excludes SB contributions dominated by dynamical diffraction. Figure 4 shows the integrated patches under both strong and weak beam conditions. Postprocessing is then applied to improve image contrast:
Background subtraction: The mean intensity plus one standard deviation is subtracted from each pixel.
Normalization: Pixel values are scaled by the maximum intensity in the image.
A fixed intensity threshold is applied to generate a binary map, highlighting potential defect regions.
Figure 4 (A) and (B) shows LCNN-labeled diffraction contrast maps under WB and SB conditions after post-processing. WB conditions enhance dislocation contrast, making dislocation boundaries and branching structures more distinct. During inference, each frame within a rocking curve is classified as WB or SB, allowing fast and consistent defect identification while excluding SB frames for efficient data reduction. Importantly, Figure 4 (C) and (D) demonstrate model predictions on previously unseen experimental data, where the LCNN generalizes successfully and correctly distinguishes WB from SB. The WB-labeled rocking layers are then assembled to reconstruct three-dimensional dislocation networks for automated and spatially resolved analysis of defect structures.
To place this work in context, it is instructive to compare our LCNN-based workflow with other relevant approaches for weak-beam condition identification in DFXM. For instance, Huang et al [25] proposed a method similar to principal component analysis (PCA) that uses Gram-Schmidt orthogonalization instead to decompose rocking curve data into three orthogonal components corresponding to the SB condition and the WB conditions on either side of the rocking curve. Unlike the data-driven approach employed here, their method does not rely on any training data but instead defines physically meaningful basis functions from manually selected regions of interest. The main advantage of such a physics-based orthogonalization is its interpretability and its grounding in the underlying diffraction contrast mechanisms. However, this also makes it more dependent on careful manual input and less straightforward to implement in an automated or real-time setting, particularly for large, heterogeneous datasets. By contrast, the lightweight convolutional neural network presented in this work enables fully automated, patch-level classification of beam conditions with minimal computational cost, allowing high-throughput processing of large volumes of DFXM data. The trade-off is that the LCNN requires representative training data and may need to be retrained when applied to new material systems with significantly different features.
Together, these two strategies illustrate complementary approaches for achieving more efficient and scalable analysis of DFXM datasets. The LCNN provides a rapid and robust solution for weak-beam identification under diverse conditions, and its model-agnostic design makes it a promising candidate for application in other imaging modalities such as TEM and X-ray topography. Moreover, we emphasize that while the model may occasionally mislabel rocking curves from unfamiliar samples, the same workflow remains valid: by manually labeling a small subset of the new dataset and retraining the model, it can be effectively adapted to the specific features of interest.
As a final validation of our approach, we reproduced the three-dimensional dislocation structures previously published in Ref. [13], using only the two-dimensional image layers automatically identified as WB conditions by our model. Figure 5 presents a direct comparison: panel (A) shows the reconstruction obtained through our fully automated workflow, while panel (B) corresponds to the manually segmented result from the original study. Despite using approximately ninety percent less data and requiring no human interpretation, the automated method yields excellent agreement with the published result, clearly demonstrating its reliability and efficiency. This capability is particularly relevant for time-resolved experiments at advanced synchrotron and free-electron laser sources, such as ESRF-EBS and XFELs, where large volumes of data are acquired rapidly.
To better understand the distinction between weak and strong beam conditions, we evaluated our model across multiple datasets. In certain cases, the separation between the two conditions is not clearly defined, making threshold-based methods unreliable. This emphasizes the advantage of our data-driven approach, which captures subtle, context-specific features beyond raw intensity. Our model offers an efficient and scalable solution for dislocation imaging across diverse materials and experimental configurations.
Conclustion and Outlook
In this work, we presented an automated and computationally efficient framework for analyzing DFXM rocking curves using lightweight convolutional neural networks. Our model enables accurate and scalable identification of weak-beam conditions crucial for interpreting diffraction contrast and visualizing dislocations in crystalline materials. This approach significantly accelerates the analysis process, opening the door to high-throughput, three-dimensional dislocation mapping in bulk samples.
Looking ahead, we aim to extend this method for real-time, on-the-fly data reduction during DFXM experiments, thereby streamlining acquisition and enabling adaptive imaging strategies. Concurrently, we are developing models to infer dislocation types and Burgers vectors by training neural networks on synthetic datasets generated from geometrical optics simulations [11, 46]. These advancements will contribute toward a more quantitative, interpretable, and fully automated framework for defect characterization in crystalline materials. The methodology presented here is also broadly applicable to other imaging modalities where weak-beam contrast plays a critical role, including X-ray topography, white-beam Laue microscopy, and transmission electron microscopy.
Methodology
Weak beam dark field microscopy
Weak Beam Dark Field (WBDF) microscopy [8] is an imaging technique used to enhance contrast from crystalline defects such as dislocations and strain fields [47, 48]. By operating slightly off the exact Bragg condition [49], WBDF suppresses background contrast and amplifies strain-induced intensity variations. This makes it particularly effective for resolving nanoscale structural distortions and has been widely applied in electron microscopy to study dislocation contrast and dynamical effects [50, 51]. When implemented with high-brilliance synchrotron sources, WBDF X-ray microscopy further enables non-destructive characterization of bulk materials, providing insight into strain distributions, phase transformations, and defect structures in metals, semiconductors, and related systems.
Data acquisition
The data used in this work originates from DFXM experiments conducted on three different samples. These experiments were performed at the beamlines ID06-HXM [52] and ID03 [15] of the European Synchrotron Radiation Facility (ESRF). The sample used in this work was a single crystal of industrially-pure aluminum (99.99%), purchased from the Surface Preparation Laboratory, with dimensions , oriented with the long axis perpendicular to the scattering plane. Prior to the experiment, the sample was annealed for 10 h at 590 ∘C in atmospheric air, then slowly cooled in the furnace. For further details see Ref. 13.
Lightweight Convolutional neural networks
Convolutional Neural Networks (CNNs) [53] are widely used deep learning models designed for image tasks such as classification, object detection, and segmentation. Unlike traditional neural networks where every neuron is connected to all others, CNNs use convolution operations to focus on small regions of an image, allowing the network to learn patterns like edges, textures, and shapes more efficiently. This structure reduces the number of parameters and improves performance.
A typical CNN is made up of several key layers:
Convolutional layers, which apply filters to extract local features from the image.
Activation functions, such as ReLU, which introduce non-linearity and help the model learn complex patterns.
Pooling layers, which reduce the size of feature maps, speeding up computations and helping prevent overfitting.
Fully connected layers, which take the learned features and output the final classification.
Standard convolutional layers use 3D filters that operate across both spatial dimensions and the full channel depth. While effective, this is computationally expensive because each filter processes all input channels simultaneously.
To improve efficiency, particularly in mobile or real-time applications, lightweight CNNs often use depthwise separable convolution, which factorizes a standard convolution into two steps:
1.
Depthwise convolution , applies a spatial filter to each input channel independently, capturing spatial patterns without mixing channels.
2.
Pointwise convolution , applies a convolution across channels to combine their information and learn inter-channel dependencies.
This separation reduces computational cost by avoiding large filters that span all channels at once. Further, a standard convolution with kernel size , input channels, and output channels, the number of parameters is:
In a depthwise separable convolution, the parameter count becomes:
where the first term is from the depthwise convolution and the second from the pointwise convolution. This factorization greatly reduces parameters and computation.
CNNs are usually trained using the backpropagation algorithm[54], which iteratively adjusts network weights to minimize classification errors. The training process involves two main steps: forward propagation and backward propagation. In forward propagation, input data passes through the network, generating predictions. The loss function then quantifies the discrepancy between predicted and true labels.
One commonly used loss function for classification tasks is the cross-entropy loss, defined as:
where is the number of training samples, is the number of classes, is a binary indicator that equals 1 if the true class for sample is , and is the predicted probability for class . The cross-entropy loss penalizes incorrect predictions, particularly when the model assigns low probability to the correct class. This makes the model produce probability distributions that align closely with true labels.
During backpropagation, the gradient of the loss function is computed with respect to each weight in the network using the chain rule of differentiation. These gradients are then utilized by optimization algorithms such as Stochastic Gradient Descent (SGD) [55] or Adaptive Moment Estimation (Adam)[56] to update the weights, minimizing the loss iteratively over multiple training epochs.
Data Availability
The codes for this study are available and accessible via the link https://github.com/Benhadjira/DL-WBDF
Author Contributions Statement
A.B. developed the machine learning model, performed the formal analysis, and processed the data. The research idea was conceived by C.D., S.B., V.F.N., and C.Y. Experimental measurements were carried out by C.Y. The initial manuscript draft was written by A.B. with input from C.Y. All authors reviewed and approved the final version of the manuscript.
Competing Interests Statement
The authors declare no competing interests.
Acknowledgement
We thank the ESRF for providing beamtime at ID06-HXM and ID03. C.Y. acknowledges financial support from the ERC Starting Grant “D-REX” No. 101116911. S.B. acknowledges support from ERC Advanced Grant nr. 885022 and from the Danish ESS lighthouse on hard materials in 3D, SOLID. We also thank Prof. L. Dresselhaus-Marais for providing the samples and for the fruitful discussions.
Funding information
This project has been partly funded by the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie COFUND scheme with grant agreement No. 101034267 and the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No. 10116911)